a motorist traveling at 72 kilometre-per-hour on citing a stop road signs apply the brakes such that under constant deceleration the car is brought to stop within a distance of 50m.what is magnitude of the deceleration



1
Expert's answer
2021-10-26T17:18:26-0400

"v=\sqrt{2|a|s},\implies"

"|a|=\frac{v^2}{2s},"

"a<0,"

"a=-\frac{20^2}{2\cdot 50}=-4~\frac m{s^2}."